I'm trying to mix oils. The recipe calls for 5,670 grams of myrrh and a hin or 1 liter of olive oil. I want to make a smaller version. i am starting with 10 ml of myrrh and need to know how much olive oil to use. I need to know how to figure out the ratio of olive oil to myrrh from the original recipe so I know how much oil I need in my smaller mixture. Please help.
thanks

Answer by Alan3354(69443)

I'm trying to mix oils. The recipe calls for 5,670 grams of myrrh and a hin or 1 liter of olive oil. I want to make a smaller version. i am starting with 10 ml of myrrh and need to know how much olive oil to use. I need to know how to figure out the ratio of olive oil to myrrh from the original recipe so I know how much oil I need in my smaller mixture.
---------
V = amount of olive oil in liters.
5760/1 = 10/V
5760V = 10
V = 1/576 liters
V =~ 1.736 milliliters
PS 1 ml = 1 cc (cubic centimeter)